# Nonasymptotic Convergence Rates for Plug-and-Play Methods With MMSE Denoisers

### v4

@@ IV Experiments
We next consider PGD for the Gaussian blurring inverse problem (1), where $X$ is sampled from the MNIST data set of handwritten digits [lecun_mnist] (resolution of $28\times 28$). $A$ is a Gaussian blurring operator, computed by convolution with a $3\times 3$ Gaussian kernel with $\sigma^{2}=1$, and measurement noise $\epsilon$ being a zero-mean standard Gaussian with variance $\sigma^{2}=0.04$. The Gaussian kernel has unit mass, giving the blur an operator norm $\lVert A\rVert_{\mathrm{op}}=1$.
We use data-fidelity term $f(x)\coloneqq\tfrac{1}{2}\|Ax-y\|^{2}$ normalized to have $L<1$. We use an MMSE denoiser trained with noise $\sigma^{2}=0.04$ for the proximal step. We confirm the efficacy of this algorithm on 5 test examples in Figure 4. At this noise level, the observed blurry images have a PSNR of $\approx 15$ dB, while the reconstructions reach PSNR of $\approx 20$ dB.
In Figure 5, we plot $\operatorname{dist}(0,\widehat{\partial}F(x_{k}))$ over $50$ iterations for the first test image in Figure 4, showing clear sublinear fixed-point convergence, as predicted by Theorem III.7.
Figure 4: PNP-PGD results on the MNIST dataset under Gaussian blur.
Figure 5: Stationary residual (41) over 50 iterations for PNP-PGD using an MMSE denoiser on the MNIST dataset under Gaussian blur.
IV-C Computed Tomography.
We study the CT inverse problem with forward operator $A$ given by the discretized Radon transform, normalized to $\|A\|_{\mathrm{op}}<1$. Training images are random $128\times 128$ patches sampled from the MayoCT dataset [mcollough_low] (resolution of $512\times 512$).
We test the PGD algorithm using the MMSE denoiser on randomly selected regions of the Shepp–Logan phantom [shepp_fourier], which captures a wide range of structures relevant in computed tomography. We corrupt measurements with zero-mean Gaussian noise to yield input PSNR levels of $20$, $24$, and $28$ dB, ranges commonly studied for low-dose CT (LDCT). See for example [zhao_low, Table 6]. After $5$ iterations of PGD we obtain reconstructions in the range of 37373838 dB which we display in Figs. 6 and 7.
Figure 6: Randomly selected $128\times 128$ region of the Shepp–Logan phantom. Clean, noisy/blurred, and PnP-PGD reconstruction with an MMSE denoiser after 5 iterations. Gaussian noise; per-panel PSNR/SSIM are denoted.
Figure 7: Randomly selected $128\times 128$ region of the Shepp–Logan phantom. Clean, noisy/blurred, and PnP-PGD reconstruction with an MMSE denoiser after 5 iterations. Gaussian noise; per-panel PSNR/SSIM are denoted.
In Fig. 8, we plot the stationary residual (41) for 160 iterations and observe behavior consistent with our findings in Theorem III.7
Figure 8: Stationary residual (41) over 160 iterations for PNP-PGD on the MayoCT dataset using an MMSE denoiser.
IV-D Architectural Details.
We build upon the implementation of the Input Convex Neural Network (ICNN) in [fang_whats] to approximate the MMSE denoiser. The authors produced excellent results with softplus activations and a mixture of linear and convolutional layers. [^4] The exact PyTorch implementation is available at https://github.com/ZhenghanFang/learned-proximal-networks The network uses layers of the form
$h_{0}$ $=\sigma(A_{0}x+b_{0}),$ (53)
$h_{k+1}$ $=\sigma(W_{k+1}h_{k}+A_{k+1}x+b_{k+1}),$ (54)
with hidden layers $W_{k}\in\mathbb{R}^{H\times H}$, skip connections $A_{k}\in\mathbb{R}^{H\times H}$, and softplus activation $\sigma(x)=\tfrac{1}{\beta}\log\mathopen{}\mathclose{{\left(1+\exp(\beta x)}}\right)$ applied element-wise ($\beta=10$). Softplus is a differentiable approximation of ReLU, so that the denoiser can be defined as the gradient $\psi_{\theta}\coloneqq\nabla g_{\theta}$ of potential $g_{\theta}:\mathbb{R}^{n}\to\mathbb{R}$ given by
$g_{\theta}(x)\coloneqq w_{\mathrm{out}}^{T}h_{L}+a_{\mathrm{out}}^{T}x+b_{\mathrm{out}}.$ (55)
To ensure convexity in $x$, all hidden layers $W_{k}$ are clipped so that they have only non-negative entries. Note that a convolutional layer can be represented as a matrix. For the MayoCT and MNIST experiments, the hidden layers are clipped in this form, not in kernel form. Because the activation is convex and non-decreasing, these constraints make $g_{\theta}$ convex, guaranteeing that its gradient ($\psi_{\theta}$) is a proximal operator [gribonval_characterization, Corollary 1]. Note that no restrictions on the skip-connections or biases are necessary.
IV-E Training Details.
For the illustration of Theorem III.2, the data was divided into batches of $4,000$ with $2,000$ iterations at learning rate $10^{-3}$. For Gaussian deblurring, the dataset was divided into batches of 500 with noise of $\sigma^{2}=0.04$. The training schedule is detailed in Table I.
TABLE I: MNIST and MayoCT Denoiser Learning-rate Schedule
Phase Learning rate Steps
1 $1\times 10^{-3}$ 10,000
2 $3\times 10^{-4}$ 15,000
3 $1\times 10^{-4}$ 10,000
4 $3\times 10^{-5}$ 5,000
For computed tomography, the dataset was divided into batches of 64 with noise level of $\sigma^{2}=0.01$. The forward operator was implemented using the ODL [^5] Documentation available at https://odlgroup.github.io/odl/. Python library and normalized to have operator norm $\|A\|_{op}<1$. Training examples were identical to Table I. All training and testing was done on an Nvidia 4070 GPU using the Ubuntu 24.04 Windows Subsystem for Linux.
Remark IV.1.
We note that while the restriction of $L<1$ in Theorem III.7 may seem potentially restrictive, any $L$-smooth data-fidelity term can be scaled appropriately to meet this condition. For example, in our computed tomography experiment, the forward operator, a discretized Radon transform is $L$-smooth with $L\approx 450$. We simply scale the data fidelity term appropriately and still obtain strong empirical results with the desired convergence rate.
@@ V Conclusion
We presented a novel representation of the MMSE denoiser’s implicit regularizer in Theorem III.2 and used it to derive nonasymptotic convergence results for PnP-PGD using an MMSE denoiser in Theorem III.7. Our analysis is strictly tied to the Gaussian-noise MMSE setting. Tweedie’s formula, for example, does not transfer to other noise models, so we suspect this flavor of result is unique to the Gaussian noise setting. However, we expect analogous nonasymptotic guarantees to be attainable for other PnP schemes that use MMSE denoisers under Gaussian noise, such as PnP-ADMM or the proximal point method. Exploration of these extensions constitutes future work toward more predictable PnP solvers.

### v6

@@ IV-C Gaussian Deblurring
We next consider the Gaussian deblurring inverse problem Eq. 1, where $x$ is sampled from the MNIST data set of handwritten digits [40], uniformly at random. The forward model $A$ is a Gaussian blurring operator, computed by circular convolution with a $(3\times 3)$ Gaussian kernel with $\sigma^{2}=1$. The kernel is normalized to have unit mass. We normalize the blur operator to have operator norm <1<1 to satisfy the assumptions in Theorem III.3.
We use the DNN architecture described in Section IV-B, trained using the Adam optimizer with a cosine annealing learning rate schedule. The network is trained with batches of $500$ for 50 00050\,000 epochs with a final learning rate of $3\times 10^{-5}$. We let $y=Ax+\epsilon$, with $\epsilon$ zero-mean Gaussian noise with variance $\sigma^{2}=0.04$. We use the data-fidelity term $f(x)\coloneq\tfrac{1}{2}\|Ax-y\|^{2}$ to get iterates of the form Eq. 7. We plot $\min_{i\leq K}\|\nabla F(x_{i})\|$ for the first $100$ iterations in Fig. 3, which we computed via Eq. 40. We display the reconstruction results in Fig. 4. Observe that there is clear sublinear decay, consistent with our theoretical results in Theorem III.3 and Section III-B.
IV-D Computed Tomography
We now consider the computed tomography inverse problem. The forward operator $A$ is given by a parallel beam Radon transform discretized via the Operator Discretization Library (ODL) [1]. We scale the operation by $1/(1+\|A\|_{\mathrm{op}})$ in accordance with Theorem III.3. We use 200 projection angles, uniformly spaced over $0,\pi)$, with 400 detector bins, implemented using the Astra toolbox [63] with CUDA acceleration.
We use the DNN architecture described in Section IV-B, trained using the Adam optimizer with a cosine annealing learning rate schedule. The network is trained with batches of $64$ for 40 00040\,000 epochs with a final learning rate of $10^{-6}$. We use training and testing data from the MayoCT dataset [44] (resolution of $512\times 512$, pixel values ranging from 0 to 1). We let $y=Ax+\epsilon$, with $\epsilon$ zero-mean Gaussian noise with variance $\sigma^{2}=0.04$. We use the data-fidelity term $f(x)\coloneq\tfrac{1}{2}\|Ax-y\|^{2}$ to get iterates of the form Eq. 7. We plot $\min_{i\leq K}\|\nabla F(x_{i})\|$ for the first $1000$ iterations in Fig. 5, which we computed via Eq. 40. We display the reconstruction results for three slices of the MayoCT dataset in Fig. 6. Observe that there is clear sublinear decay, consistent with our theoretical result in Theorem III.3 and Section III-B.
@@ V Conclusion
We presented a novel representation of the MMSE denoiser’s implicit regularizer in Theorem III.2 and used it to derive nonasymptotic convergence results for PnP-PGD using an MMSE denoiser in Theorem III.3 and Section III-B. Our analysis is strictly tied to the Gaussian-noise MMSE setting. Tweedie’s formula, for example, is not the same under other noise models [22], so we suspect this flavor of result is unique to the Gaussian setting. However, we expect analogous nonasymptotic guarantees to be attainable for other PnP schemes that use MMSE denoisers under Gaussian noise, such as PnP-ADMM or the proximal point method. Exploration of these extensions constitutes future work.